# LetsGo: Large-Scale Garage Modeling and Rendering via LiDAR-Assisted Gaussian Primitives

### v2

@@ 3.2 GarageWorld Dataset
We collect data for eight garagesAll garage data is captured with the necessary permits., including six underground garages with various types of inside geometries, one indoor garage with multi-floors at Shopping Mall One, and one outdoor surface parking at Shopping Mall Two. We plan to share the datasets with the community for further research. These garages comprise various challenging structures, such as sloped surfaces with distorted lines, internal circular or spiral paths, vehicle elevators, etc., as shown in Fig. 5. Tab. 1 provides an overall description of the garages.
@@
3.3 Initial Mesh Reconstruction
We employ the off-the-shell LiDAR-Inertial-Visual(LIV) SLAM [69] to estimate the relative poses of the sensor between different time steps. The LIV-SLAM system integrates a tightly coupled LiDAR-Inertial Odometry and Visual-Inertial Odometry, along with a joint optimization approach between LiDAR and camera data, for relative pose estimation. The pose estimation system leverages the unique capabilities of each sensor of our Polar device to enhance overall accuracy and robustness. The IMU delivers reliable short-term motion estimates, while the LiDAR contributes precise distance measurements. Visual sensors complement these by enriching pose estimation in environments abundant with visual features, ensuring a robust and precise outcome.
After merging the point cloud data collected at different time steps using the estimated relative pose, we apply Poisson Reconstruction [32] to convert the point cloud data into a mesh. To ensure the accuracy and integrity of the resulting mesh, we also compare the reconstructed meth to the original point cloud data to remove incorrect faces (More details in Appendix.B). This geometry-based mesh reconstruction method performs well for Lambertian surface reconstruction. However, it faces challenges for reflective and transparent surfaces, e.g., vehicle glass windows, which is especially common in garage environments.
4 LiDAR-Assisted LoD Gaussian
3D Gaussian representation excels at modeling transparent and reflective surfaces compared to mesh. However, it requires a sparse 3D point cloud of the scene and accurate camera parameters obtained from SfM. As discussed previously, the large-scale garage scenes are challenging for SfM algorithms, with low lighting conditions, large texture-less regions, repetitive patterns, etc., making sufficient feature correspondences between different images hard to establish. As a result, the capacity of the original 3DGS for high-quality modeling and rendering is limited.
@@ 4 LiDAR-Assisted LoD Gaussian
4.1 Gaussian Splatting with LiDAR Inputs
3DGS [34] represents 3D points with 3D Guassians, parameterized by position $\mu$, opacity $\alpha$, anisotropic covariance $\Sigma$, and spherical harmonic (SH) coefficients representing view-dependent color $c$. The projection from 3D Gaussians to 2D images [101] is given by
$\Sigma^{\prime}=JW\Sigma W^{T}J^{T},$ (1)
@@ 4.1 Gaussian Splatting with LiDAR Inputs
The purpose is to optimize the Gaussian parameters so that the rendered images from the 3D Gaussians are as close to their ground truth (GT) images as possible. Normally, the optimization is achieved by stochastic gradient descent (SGD). Since it cannot constrain the covariance matrix to be semi-definite, only in which scenario the covariance matrix has its physical meaning, Kerbl et al. proposes to optimize a scaling matrix $S$ and rotation matrix $R$, and compute the covariance matrix as:
$\Sigma=RSS^{T}R^{T},$ (2)
where $S$ is parameterized as a 3D vector and $R$ is parameterized as quaternion, a 4D vector with an unit norm.
Since conventional SfM approaches fail in large garage environments due to hard correspondence establishment between textureless and transparent regions, we use our Polar device with a LiDAR sensor to scan point clouds of the garages. Considering the originally scanned point clouds contain noises, we resample a set of new points from the reconstructed mesh with a uniform sampling strategy. These resampled points, in conjunction with the camera parameters, are used to train the 3DGS representations.
@@
Depth Regularizer. Inspired by the depth calculation from NeRF [51], we utilize the rasterization pipeline of Gaussians to compute the depth of each Gaussian primitive:
$D_{\text{G}}=\sum_{i\in N}d_{i}\alpha_{i}T_{i},\quad\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$ (3)
@@
$d_{i}=\frac{1}{l}\left(p_{2}-\frac{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{0,2}}{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{2,2}}\left(x_{0}-p_{0}\right)-\frac{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{1,2}}{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{2,2}}\left(x_{1}-p_{1}\right)\right)$ (4)
@@
$\mathcal{L}_{\mathrm{depth}}=\sum_{k}\left\|{D}^{k}_{G}-D^{k}\right\|_{1},$ (5)
where $D^{k}$ represents the inherent depth prior from LiDAR data.
@@
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{rgb}}+\lambda_{\mathrm{depth}}\mathcal{L}_{\mathrm{depth}},$ (6)
@@
Figure 4: Illumination of our LOD-LiDAR-GS within an Octree Structure, accompanied by an LOD rendering strategy. Distance varies from near to far, with Gaussian level resolution transitioning from fine to coarse. The bottom row illustrates the level composition in the rendered view.
4.2 Multi-Resolution Representation
@@
Inspired by advanced rendering techniques [66] for massive LiDAR point clouds, we introduce a multi-resolution Gaussian framework called LOD-LiDAR-GS. This framework integrates the Level-of-Detail (LOD) rendering to the LiDAR-GS, making it suitable for various devices. We represent the 3D scene with Gaussians at different resolution levels. Each Gaussian has an additional LOD value, alongside its original attributes like 3D position, spherical harmonic coefficients, and opacity. The LOD value dynamically adjusts the cloning and splitting threshold during training. This multi-resolution representation enables fast and lightweight rendering on web-based devices.
Figure 5: Visualization of examples from our GarageWorld dataset.
@@
We first construct multi-resolution point cloud data for Gaussian initialization. Our Polar device generates dense point clouds with superior accuracy and consistency, even in challenging garage conditions. We use a spacing attribute, $\tau$, to represent the resolution of a point cloud, defined as the minimal distance between points. The finest resolution corresponds to the original point cloud sampled from our reconstructed mesh at the lowest $\tau$. We then increase $\tau$ to $2\tau$ and downsample the point cloud to derive a coarser point cloud, repeating until the point count drops below a threshold, $\epsilon_{p}$. Points in the finest level get an LOD value of $N=L-1$, and those in the coarsest level get an LOD value of $N=0$. We set $\tau$ to 4.0 cm and $\epsilon_{p}$ to 10,000 points. The downsampling operation employs an approximate Poisson-disk sampling method from the PotreeConverter library [66].
LOD-LiDAR-GS Training.
In contrast to LiDAR-GS, our LOD-LiDAR-GS framework processes a multi-resolution point cloud, wherein each resolution level has a distinct LOD value. We devise a new training strategy to handle the complexity caused by multi-resolution input effectively. We initialize a Gaussian model for each resolution level, which operates independently during optimization. Following the original 3DGS [34], the decision to clone or split a Gaussian is based on threshold values for position gradients $\sigma_{pos}$ and variance $\sigma_{var}$. We introduce a scaling factor, $s$, which varies with the LOD value $k$ of each level, modifying the thresholds as $s\cdot\sigma_{pos}$ and $s\cdot\sigma_{var}$:
$s_{k}=\min(\beta_{s}^{L-1-k},s_{max}),\quad k\in[0,L-1]$ (7)
$L^{(d)}=\text{clamp}(\lfloor L^{1-d/d_{max}}\rfloor,0,L-1)$ (8)
@@
4.3 Web-based Lightweight Renderer
With the proposed multi-resolution Gaussian representation, our LOD-LiDAR-GS significantly enhances rendering efficiency. Instead of using a large number of 3D Gaussians for each image, LOD-LiDAR-GS dynamically selects Gaussian subsets from different resolution levels based on depth ranges. This reduces the complexity of sorting, projection, and accumulation operations, thereby accelerating rendering speed. On GPU systems such as the NVIDIA RTX3090, we implement a 3D Gaussian viewer based on the original 3DGS SIBR viewer. By integrating our LOD rendering strategy, as detailed in Equation 8, we achieve approximately a 4x increase in rendering speed. Additionally, we introduce a lightweight renderer employing an on-demand rendering approach. Unlike mainstream lightweight Gaussian renderers [39, 18], which load all Gaussians into VRAM simultaneously, our renderer dynamically loads only the necessary data into memory according to the current viewpoint and LOD strategy. This method avoids the need for full data loading, enabling the rendering of large-scale environments on web platforms.
Figure 6: Illustration of our partition scheme. The left image shows the margin expansion method employed to guarantee overlap amongst the partitioned blocks. The top right image demonstrates the extension of partition boundaries to incorporate irregular areas. The bottom right image details the approach for multi-level parking structures, wherein each level is segregated into different partitions.
@@ 4.3 Web-based Lightweight Renderer
Rendering large-scale garage scenes requires managing a substantial number of Gaussians, often exceeding in-memory storage capacity. To address this, we implement a coarse-to-fine loading and rendering scheme that facilitates real-time rendering on resource-constrained devices. Initially, low-resolution Gaussian levels are preloaded into memory at viewer initialization. These levels occupy minimal storage and are accessed frequently during subsequent rendering cycles. As shown in Fig. 4, to render high-frequency details, our system dynamically selects high-resolution octree chunks from disk based on the viewer’s frustum and the conditions set by our LOD rendering strategy, as outlined in Equation 8. The rendering process begins with a traversal of the entire octree, determining each node’s visibility based on the center position of the node’s bounding box. A second traversal then loads visible Gaussian data into a buffer, continuing until either the traversal is complete or the number of loaded Gaussians reaches the maximum capacity. The buffer is updated by replacing preloaded Gaussians with newly loaded high-resolution Gaussians, which are then dispatched to a WebGL worker for rendering. Recent works, such as OCT-GS [61] and LoG [70], can also be structured into an octree format. However, OCT-GS fails to represent the entire scene at certain levels, making it unsuitable for our lightweight rendering strategy. Additionally, LoG does not provide a solution for lightweight rendering.
5 Experimental Results
5.1 Training Details
Figure 7: Qualitative comparison between our LOD-LiDAR-GS, 3DGS, Mip-Splatting, F2-NeRF, NGP and Splatfacto on the various datasets. Here, we use a superscription to denote their Gaussian primitives are initialized according to our LiDAR-assisted 3D point cloud.
Table 2: Quantitative comparison among our LOD-LiDAR-GS, 3DGS, Mip-Splatting, F2-NeRF, NGP, Splatfacto, OCT-GS, and LoG-GS on the various datasets.
@@ 5.1 Training Details
All experiments are conducted on an NVIDIA RTX A6000 GPU using the PyTorch Framework. Given the large scale of our scenes and the initialization with LiDAR point clouds, we set the scaling learning rate to 0.0015 and the initial position learning rate for Gaussian primitives to 0.000016. We disable the opacity reset option and delay the start of the Gaussian densification step to 75,000 iterations. The total number of iterations is empirically set to twenty times the number of captured images. We use a spherical harmonics (SH) degree of 2 and set $\lambda_{\mathrm{depth}}$ to 0.8.
Block Partitioning and Merging.
To accelerate our model training through parallel execution, we employ a partition-based strategy. As illustrated in Fig. 6, each garage floor is divided into sub-blocks using overlapping quadrilaterals, typically rectangles or trapezoids with small internal angles. Each block covers approximately 100 square meters, enabling efficient small-scale training for scene representation. To address the poor reconstruction quality at the edges, common in Gaussian splatting-based methods, we expand each quadrilateral outward by 30%. This expansion enhances the overall reconstruction quality. After training each partition, we merge the results using the original quadrilateral coordinates, achieving a complete scene reconstruction. This method ensures high-quality rendering while optimizing the training process.
@@
5.2 Comparison
@@
We compare our approach with state-of-the-art methods and concurrent works on various challenging datasets, including our GarageWorld, KITTI-360 [44] (large outdoor street scene), and ScanNet++ [92] (indoor scene characterized by complex geometry and variable lighting). For each dataset, we use 10% of the images as test sets and the remaining 90% as training sets.
@@
We compare our method against a suite of 3DGS approaches, including 3DGS* [34], Mip-Splatting* [94], and Splactfacto* [76], as well as recent concurrent work, including OCT-GS* [61] and LoG* [70]. Additionally, we benchmark our approach against recent implicit representation methods, namely F2-NeRF [81] and Instant-NGP [53]. To extend recent 3DGS-based methods to large-scale garage scenes, we initialize them with our LiDAR-derived point clouds instead of SFM point clouds, enabling their application to challenging garage environments. Methods utilizing our LiDAR-derived point clouds are denoted with an asterisk (*).
Figure 8: Qualitative comparison of our method and concurrent works, including recent OCT-GS* and LoG* methods, on the GarageWorld and UAVD4L datasets.
Qauantitative Comparison Results.
For quantitative comparisons, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS)[96] as metrics to evaluate rendering quality. The results of these comparisons are presented in Tab. 2. Our LOD-LiDAR-GS method consistently achieves the best results on large-scale scene datasets, such as our GarageWorld dataset and the KITTI-360 dataset. On small-scale scene dataset ScanNet++, our LOD-LiDAR-GS approach achieves the second-best results. From the last two rows of Tab. 2, our web renderer demonstrate comparable performance on lightweight devices (MacBook Air with M2 clip).
Figure 9: Qualitative ablation study on depth regularization within our large-scale GarageWorld dataset clearly demonstrates improvements in resolving the floater problem on the ground.
Figure 10: Ablation study on our LOD rendering strategy demonstrates that our real-time LOD renderer operates approximately 3-4 times faster than the original 3DGS SIBR viewer when tested on an NVIDIA RTX 3090 GPU.
@@
As shown in Fig. 7, both 3DGS* and Splactfacto* tend to generate floating Gaussian points around scene surfaces, resulting in blurriness and poor rendering quality. F2-NeRF performs well on the relatively small-scale ScanNet++ dataset but suffers from artifacts in the large-scale KITTI-360 dataset and blurriness in our GarageWorld dataset. Instant-NGP, which utilizes hash encoding, produces notably blurry results in large-scale scenes. In contrast, our method demonstrates the highest rendering quality across various scenarios. We also compared our method with concurrent works such as OCT-GS* and LoG*. As shown in Fig. 8, our approach achieves the best rendering quality on both our GarageWorld and the UAVD4L datasets (recently proposed and used in LoG*). Our web-based results also show comparable visual quality.
@@
Additionally, we evaluate the performance of our web-based renderer on various devices. On a high-performance desktop equipped with an i9-10900X CPU and a Samsung SSD T5 Disk, it takes 1.36 seconds to load approximately 2 million 3D Gaussians. On a MacBook laptop with an Apple M2 CPU and Apple SSD AP0512Z Disk, the load time for the same number of 3D Gaussians is 1.29 seconds. Combining these newly loaded 3D Gaussians with an additional 3 million pre-loaded Gaussians, both devices maintain a rendering frame rate of 60 FPS.
5.3 Ablation Study

### v3

@@ 3.2. GarageWorld Dataset
We collect data for eight garages [^1] All garage data is captured with the necessary permits., including six underground garages with various types of inside geometries, one indoor garage with multi-floors at Shopping Mall One, and one outdoor surface parking at Shopping Mall Two. These garages comprise various challenging structures, such as sloped surfaces with distorted lines, internal circular or spiral paths, vehicle elevators, etc., as shown in Fig. 4. Tab. 1 provides an overall description of the garages.
@@
3.3. Initial Mesh Reconstruction
We employ the off-the-shell LiDAR-Inertial-Visual (LIV) SLAM (Shan et al., 2021) to estimate the relative poses of the sensor between different time steps. The LIV-SLAM system integrates a tightly coupled LiDAR-Inertial Odometry and Visual-Inertial Odometry, along with a joint optimization approach between LiDAR and camera data, for relative pose estimation. The pose estimation system leverages the unique capabilities of each sensor of our Polar device to enhance overall accuracy and robustness. The IMU delivers reliable short-term motion estimates, while the LiDAR contributes precise distance measurements. Visual sensors complement these by enriching pose estimation in environments abundant with visual features, ensuring a robust and precise outcome.
After merging the point cloud data collected at different time steps using the estimated relative pose, we apply Poisson Reconstruction (Kazhdan et al., 2006) to convert the point cloud data into a mesh. To ensure the accuracy and integrity of the resulting mesh, we also compare the reconstructed mesh to the original point cloud data to remove incorrect faces (More details in Appendix. B). This geometry-based mesh reconstruction method performs well for Lambertian surface reconstruction. However, it faces challenges for reflective and transparent surfaces, e.g., vehicle glass windows, which is especially common in garage environments.
4. LiDAR-Assisted LoD Gaussian
3D Gaussian representation excels at modeling transparent and reflective surfaces compared to a mesh representation. However, it requires a sparse 3D point cloud of the scene and accurate camera parameters obtained from SfM. As discussed previously, the large-scale garage scenes are challenging for SfM algorithms, with low lighting conditions, large texture-less regions, repetitive patterns, etc., making sufficient feature correspondences between different images hard to establish. As a result, the capacity of the original 3DGS for high-quality modeling and rendering is limited.
@@ 4. LiDAR-Assisted LoD Gaussian
4.1. Gaussian Splatting with LiDAR Inputs
3DGS (Kerbl et al., 2023) represents 3D points with 3D Gaussians, parameterized by position $\mu$, opacity $\alpha$, anisotropic covariance $\Sigma$, and spherical harmonic (SH) coefficients representing view-dependent color $c$. The projection from 3D Gaussians to 2D images (Zwicker et al., 2001b) is given by
(1) $\Sigma^{\prime}=JW\Sigma W^{T}J^{T},$
@@ 4.1. Gaussian Splatting with LiDAR Inputs
The goal is to optimize the Gaussian parameters so that the rendered images from the 3D Gaussians are as close to their ground truth (GT) images as possible.
A standard three-dimensional Gaussian sphere can be transformed into an ellipsoid of arbitrary shape within space by applying a covariance matrix $\Sigma$. Additionally, so as to ensure the semi-definiteness of the covariance matrix during the optimization process, Kerbl et al. proposes to optimize a scaling matrix $S$ and rotation matrix $R$, and compute the covariance matrix as:
(2) $\Sigma=RSS^{T}R^{T},$
where $S$ is parameterized as a 3D vector and $R$ is parameterized as quaternion, a 4D vector with unit norm.
Since conventional SfM approaches fail in large garage environments due to hard correspondence establishment between textureless and transparent regions, we use our Polar device with a LiDAR sensor to scan point clouds of the garages. Considering the originally scanned point clouds contain noise, we resample a set of new points from the reconstructed mesh with a uniform sampling strategy. These resampled points, in conjunction with the camera parameters, are used to train the 3DGS representations.
@@
Depth Regularizer. Inspired by the depth calculation from NeRF (Mildenhall et al., 2021), we utilize the rasterization pipeline of Gaussians to compute the depth of each Gaussian primitive:
(3) $D_{\text{G}}=\sum_{i\in N}d_{i}\alpha_{i}T_{i},\quad\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$
@@
(4) $d_{i}=\frac{1}{l}\left(p_{2}-\frac{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{0,2}}{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{2,2}}\left(x_{0}-p_{0}\right)-\frac{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{1,2}}{\left(\boldsymbol{\mathbf{\Sigma}}^{-1}\right)_{2,2}}\left(x_{1}-p_{1}\right)\right)$
@@
(5) $\mathcal{L}_{\mathrm{depth}}=\sum_{k\in K}\left\|{D}^{k}_{G}-D^{k}\right\|_{1},$
where $D^{k}$ represents the ${k^{\text{th}}}$ inherent depth prior from LiDAR data.
@@
(6) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{rgb}}+\lambda_{\mathrm{depth}}\mathcal{L}_{\mathrm{depth}},$
@@
4.2. Multi-Resolution Representation
@@
Inspired by advanced rendering techniques (Schütz et al., 2016) for massive LiDAR point clouds, we introduce a multi-resolution Gaussian framework called LOD-LiDAR-GS. This framework integrates the Level-of-Detail (LOD) rendering to the LiDAR-GS, making it suitable for various devices. We represent the 3D scene with Gaussians at different resolution levels. When initializing the multi-resolution Gaussians, each Gaussian is assigned an LOD attribute, alongside its original attributes like 3D position, spherical harmonic coefficients, and opacity, to support our LOD rendering. The LOD value dynamically adjusts the cloning and splitting threshold during training. This multi-resolution representation enables fast and lightweight rendering on web-based devices.
Figure 4. Visualization of examples from our GarageWorld dataset.
@@
We first construct multi-resolution point cloud data for Gaussian initialization. Our Polar device generates dense point clouds with superior accuracy and consistency, even in challenging garage conditions. We use a spacing attribute, $\tau$, to represent the resolution of a point cloud, defined as the minimal distance between points. The finest resolution corresponds to the original point cloud sampled from our reconstructed mesh at the lowest $\tau$. We then increase $\tau$ to $2\tau$ and downsample the point cloud to derive a coarser point cloud, repeating until the point count drops below a threshold, $\epsilon_{p}$. Points in the finest level get an LOD value of $N=L-1$, and those in the coarsest level get an LOD value of $N=0$. We set $\tau$ to 4.0 cm and $\epsilon_{p}$ to 10,000 points. The downsampling operation employs an approximate Poisson-disk sampling method from the PotreeConverter library (Schütz et al., 2016).
Figure 5. Illustration of our LOD-LiDAR-GS within an Octree Structure, accompanied by an LOD rendering strategy. Distance varies from near to far, with Gaussian level resolution transitioning from fine to coarse. The bottom row illustrates the level composition in the rendered view.
@@
LOD-LiDAR-GS Training.
In contrast to LiDAR-GS, our LOD-LiDAR-GS framework processes a multi-resolution point cloud, wherein each resolution level has a distinct LOD value. We devise a new training strategy to handle the complexity caused by multi-resolution input effectively. We initialize a Gaussian model for each resolution level, which operates independently during optimization. Following the original 3DGS (Kerbl et al., 2023), the decision to clone or split a Gaussian is based on threshold values for position gradients $\sigma_{pos}$ and variance $\sigma_{var}$. We introduce a scaling factor $s$, which varies with the LOD value $l$ of each level, modifying the thresholds as $s\cdot\sigma_{pos}$ and $s\cdot\sigma_{var}$:
(7) $s_{k}=\min(\beta_{s}^{L-1-l},s_{max}),\quad l\in[0,L-1]$
(8) $L^{(d)}=\text{clamp}(\lfloor L^{1-d/d_{max}}\rfloor,0,L-1)$
4.3. Web-based Lightweight Renderer
With the proposed multi-resolution Gaussian representation, our LOD-LiDAR-GS significantly enhances rendering efficiency. Instead of using a large number of 3D Gaussians for each image, LOD-LiDAR-GS dynamically selects Gaussian subsets from different resolution levels based on depth ranges. This reduces the complexity of sorting, projection, and accumulation operations, thereby accelerating rendering speed. On GPU systems such as the NVIDIA RTX 3090, we implement a 3D Gaussian viewer based on the original 3DGS SIBR viewer. By integrating our LOD rendering strategy, as detailed in Equation 8, we achieve approximately a 4x increase in rendering speed. Additionally, we introduce a web-based, on-demand renderer developed with JavaScript and WebGL, which supports real-time rendering on lightweight devices. Unlike mainstream lightweight Gaussian renderers (Kwok, 2023; Face, 2024), which load all Gaussians into VRAM simultaneously, our renderer dynamically loads only the necessary data into memory according to the current viewpoint and LOD strategy. This method avoids the need for full data loading, enabling the rendering of large-scale environments on web platforms.
@@ 4.3. Web-based Lightweight Renderer
Rendering large-scale garage scenes requires managing a substantial number of Gaussians, often exceeding in-memory storage capacity. To address this, we implement a coarse-to-fine loading and rendering scheme that facilitates real-time rendering on resource-constrained devices. Initially, low-resolution Gaussian levels are preloaded into memory at viewer initialization. These levels occupy minimal storage and are accessed frequently during subsequent rendering cycles. As shown in Fig. 5, to render high-frequency details, our system dynamically selects high-resolution octree chunks from disk based on the viewer’s frustum and the conditions set by our LOD rendering strategy, as outlined in Equation 8. The rendering process begins with a traversal of the entire octree, determining each node’s visibility based on the center position of the node’s bounding box. A second traversal then loads visible Gaussian data into a buffer, continuing until either the traversal is complete or the number of loaded Gaussians reaches the maximum capacity. The buffer is updated by replacing preloaded Gaussians with newly loaded high-resolution Gaussians, which are then dispatched to a WebGL worker for rendering. Recent works, such as OCT-GS (Ren et al., 2024) and LoG (Shuai et al., 2024), can also be structured into an octree format. However, OCT-GS fails to represent the entire scene at certain levels, making it unsuitable for our lightweight rendering strategy. Additionally, LoG does not provide a solution for lightweight rendering.
5. Experimental Results
5.1. Training Details
Figure 6. Qualitative comparison between our LOD-LiDAR-GS, 3DGS, Mip-Splatting, F2-NeRF, NGP and Splatfacto on the various datasets. Here, we use a superscript to denote their Gaussian primitives are initialized according to our LiDAR-assisted 3D point cloud.
Table 2. Quantitative comparison among our LOD-LiDAR-GS, 3DGS, Mip-Splatting, F2-NeRF, NGP, Splatfacto, OCT-GS, and LoG-GS on the various datasets.
@@ 5.1. Training Details
All experiments are conducted on an NVIDIA RTX A6000 GPU using the PyTorch Framework. Given the large scale of our scenes and the initialization with LiDAR point clouds, we set the scaling learning rate to 0.0015 and the initial position learning rate for Gaussian primitives to 0.000016. We disable the opacity reset option and delay the start of the Gaussian densification step to 75,000 iterations. The total number of iterations is empirically set to twenty times the number of captured images. We use a spherical harmonics (SH) degree of 2 and set $\lambda_{\mathrm{depth}}$ to 0.8. We employ a partitioning strategy similar to (Tancik et al., 2022), expanding each block outward by 30% to address the poor reconstruction quality on the boundaries. The overall end-to-end duration, including data capture, preprocessing, and training, typically ranges from 12 to 16 hours. Each scene is divided into 3 to 10 blocks, depending on its complexity. The training duration for each block varies from 1.5 to 4 hours, based on the size of the point cloud and the number of images. In a few cases, irregular shapes in underground garages result in particularly large blocks, extending the training time to over 8 hours.
@@
5.2. Comparison
Figure 7. Qualitative comparison of our method and concurrent works, including recent OCT-GS* and LoG* methods, on the GarageWorld and UAVD4L datasets.
Figure 8. Qualitative and Quantitative comparison of our method and READ approach on the GarageWorld datasets.
@@ 5.2. Comparison
We compare our approach with state-of-the-art methods and concurrent works on various challenging datasets, including our GarageWorld, KITTI-360 (Liao et al., 2022) (large outdoor street scene), and ScanNet++  (Yeshwanth et al., 2023) (indoor scene characterized by complex geometry and variable lighting). For each dataset, we use 10% of the images as test sets and the remaining 90% as training sets.
Figure 9. Qualitative ablation study on depth regularization within our large-scale GarageWorld dataset clearly demonstrates improvements in resolving the floater problem on the ground.
Figure 10. Ablation study on our LOD rendering strategy demonstrates that our real-time LOD renderer operates approximately 3-4 times faster than the original 3DGS SIBR viewer when tested on an NVIDIA RTX 3090 GPU.
@@
We compare our method against a suite of 3DGS approaches, including 3DGS* (Kerbl et al., 2023), Mip-Splatting* (Yu et al., 2024), and Splactfacto* (Tancik et al., 2023), as well as recent concurrent work, including OCT-GS* (Ren et al., 2024) and LoG* (Shuai et al., 2024). Additionally, we benchmark our approach against recent implicit representation methods, namely F2-NeRF (Wang et al., 2023) and Instant-NGP (Müller et al., 2022). To extend recent 3DGS-based methods to large-scale garage scenes, we initialize them with our LiDAR-derived point clouds instead of SFM point clouds, enabling their application to challenging garage environments. Methods utilizing our LiDAR-derived point clouds are denoted with an asterisk (*).
Quantitative Comparison Results.
For quantitative comparisons, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) as metrics to evaluate rendering quality. The results of these comparisons are presented in Tab. 2. Our LOD-LiDAR-GS method consistently achieves the best results on large-scale scene datasets, such as our GarageWorld dataset and the KITTI-360 dataset. On the small-scale scene dataset ScanNet++, our LOD-LiDAR-GS approach achieves the second-best results. From the last two rows of Tab. 2, our web renderer demonstrate comparable performance on lightweight devices (MacBook Air with M2 clip).
Figure 11. Qualitative ablation study on different input point cloud qualities demonstrates that using point clouds sampled from the mesh significantly enhances rendering quality.
@@
As shown in Fig. 6, both 3DGS* and Splactfacto* tend to generate floating Gaussian points around scene surfaces, resulting in blurriness and poor rendering quality. F2-NeRF performs well on the relatively small-scale ScanNet++ dataset but suffers from artifacts in the large-scale KITTI-360 dataset and blurriness in our GarageWorld dataset. Instant-NGP, which utilizes hash encoding, produces notably blurry results in large-scale scenes. In contrast, our method demonstrates the highest rendering quality across various scenarios. We also compared our method with concurrent works such as OCT-GS* and LoG*. As shown in Fig. 7, our approach achieves the best rendering quality on both our GarageWorld and the UAVD4L datasets (recently proposed and used in LoG*). Our web-based results also show comparable visual quality. Moreover, we evaluate our method against the READ method (Li et al., 2023) which also uses LiDAR as input. Fig. 8 presents the qualitative and quantitative results, showing that our method can recover more high-frequency details.
@@
Additionally, we evaluate the performance of our web-based renderer on various devices. On a high-performance desktop equipped with an i9-10900X CPU and a Samsung SSD T5 Disk, it takes 1.36 seconds to load approximately 2 million 3D Gaussians. On a MacBook laptop with an Apple M2 CPU and Apple SSD AP0512Z Disk, the load time for the same number of 3D Gaussians is 1.29 seconds. Combining these newly loaded 3D Gaussians with an additional 3 million pre-loaded Gaussians, both devices maintain a rendering frame rate of 60 FPS. Due to browser memory constraints, we restrict the 3D Gaussian data loaded for rendering to a maximum of 2 GB. Furthermore, the web renderer uses CPU-based sorting, which is slower than the CUDA-based sorting employed by the original 3DGS renderer. To balance rendering quality and performance, we adopt our LOD rendering strategy that adjusts the number of loaded 3D Gaussians.
Figure 12. Qualitative ablation study on the random-resolution-level (RRL) training strategy reveals that our RRL approach effectively addresses the overfitting issue at a certain level, influenced by the camera distribution.
5.3. Ablation Study